# Investigating cannibalistic millisecond pulsar binaries using MESA : new constraints from pulsar spin and mass evolution

### v1

@@ Abstract
Context. Compact binary millisecond pulsars (MSPs), with orbital periods $\lesssim 1$d, are key to understanding binary evolution involving massive neutron stars (NSs). Due to the ablation of the companion by the rapidly spinning pulsar, these systems are also known as spiders, categorized into two main branches: redbacks (RBs; companion mass in the range of 0.1 to 0.5 M) and black widows (BWs; companion mass $\lesssim$ 0.1 M).
Aims. We present models of low- and intermediate-mass X-ray binaries (XRBs) and compare them with observations of Galactic spiders (including the presence/absence of hydrogen lines in their optical spectra). We constrain and quantify the interaction between pulsar and companion.
Methods. Using MESA, we create the allowed initial parameter space, modeling the irradiation of the companion by the pulsar wind. For the first time in MESA , we also include the detailed evolution of the pulsar spin.
Results. Efficient mass accretion onto the NS ($70\%$ of the mass transferred) with an X-ray irradiated disc, followed by strong irradiation of the companion can explain most properties of the observed spiders. Our RB evolutionary tracks continue to the BW regime, connecting the two branches of spiders. Our models explain the lack of hydrogen in some observed BWs with ultra-light companions. The mass required to spin-up a NS to sub-milliseconds is high enough to collapse it into a black hole. Finally, we discuss the formation of RB-like spiders with giant companions and orbital periods of several days (huntsmen), concluding that they are unlikely to produce super-massive NSs (maximum accreted mass $\lesssim$0.5 ${\rm M}_{\odot}$).
@@ 1 Introduction
The formation of radio millisecond pulsars (MSPs) has long been connected to low-mass X-ray binaries (LMXBs) (Alpar et al., 1982; Radhakrishnan & Srinivasan, 1982; Bhattacharya & van den Heuvel, 1991). Observed MSPs have short spin periods ($\lesssim 30$ ms) and low surface magnetic fields ($\sim 10^{8}$ G). The neutron stars (NS) in LMXBs accrete matter which spins them up to millisecond periods (Alpar et al., 1982). This phase is observed as an X-ray phase and is also called the recycling phase due to the pulsar spin-up. Observations of accreting X-ray MSPs confirmed the link between MSPs and accreting LMXBs (Wijnands & van der Klis, 1998; Archibald et al., 2009). The NS surface magnetic field is also suggested to decrease with accreted mass due to Ohmic decay, explaining the observed magnetic field strengths of MSPs (Konar & Bhattacharya, 1997; Zhang, 1998; Konar & Bhattacharya, 1999a, b; Cumming et al., 2001). Compact binary MSPs are a sub-category of MSP binaries with compact orbits and eclipses in their radio signals. In compact binary MSPs, after the mass-transfer phase or the Roche-lobe overflow (RLO) has ended, the highly spinning pulsar loses energy in the form of energetic particles that irradiates the outer surface of the companion leading to mass loss (Archibald et al., 2009). Initial observations of eclipsing MSPs showed that some of them have very low-mass companions (0.02 to 0.05 M; Fruchter et al., 1988; Stappers et al., 1996a, 1999), which could be explained by mass loss due to pulsar wind irradiation. Evidence of irradiation-induced mass loss has since been observed in multiple eclipsing compact MSPs (Kluzniak et al., 1988; Ruderman et al., 1989; Stappers et al., 1996b; Polzin et al., 2020; Kumari et al., 2024).
@@
Due to the cannibalistic nature of this interaction, these compact binary MSPs, with observed orbital periods $P_{\rm orb}\lesssim 1\,\rm d$ (Pallanca et al., 2012; Romani et al., 2012; Kaplan et al., 2013; Breton et al., 2013; Roberts, 2013), are also called spider pulsars. So far, about 60 spiders have been observed and studied in our Galaxy (for instance see Abdo et al., 2009; Pletsch et al., 2012; Pallanca et al., 2012; Romani et al., 2012; Kaplan et al., 2013; Breton et al., 2013). There are two main categories of spiders (Roberts, 2013), based on the masses of the companions to the pulsar (denoted as $M_{\rm c}$): Redbacks (RBs; having companions with masses $0.1{\,{\rm M}_{\odot}}\lesssim M_{\rm c}\lesssim 0.5\,{\rm M}_{\odot}$) and black widows (BWs; having companions with masses $M_{\rm c}<0.1\,{\rm M}_{\odot}$). The orbital period distribution is similar for both types ($\lesssim$1 d).
@@
Following classical binary evolution, the observed orbital periods of all spiders are not explained, as they fall short by about an order of magnitude (Podsiadlowski et al., 2002). Fast winds that leave a binary isotropically and without interacting with anything else have a widening effect on the orbit (Soberman et al., 1997). Therefore, the involvement of this energetic radio pulsar wind has been invoked to explain the orbital properties of recycled pulsars (Ergma & Fedorova, 1992; Podsiadlowski et al., 2002; Benvenuto et al., 2014, 2015; Chen et al., 2013). There are a few ideas that have been proposed to explain spider MSPs. Chen et al. (2013) proposed strong and weak irradiation producing RBs and BWs, respectively. Another group studied the formation of spiders by including X-ray irradiation feedback and found that in order to reproduce the orbital parameters of spiders, cyclic mass transfer due to the effect of X-rays is required (Benvenuto et al., 2014, 2015). Ginzburg & Quataert (2020) and Ginzburg & Quataert (2021) argued that sustained magnetic braking (that couples to stellar wind leading to angular momentum loss) is needed to reproduce spiders and ablation alone is not sufficient. Some spiders are also thought to come from ultra-compact X-ray binaries (UCXBs), which are a sub-category of LMXBs with hydrogen-deficient companions and orbital periods of less than an hour (see review by Nelemans & Jonker, 2010).
@@
Since spider binaries have passed through a long mass-transfer phase, it has been suggested that they host some of the most massive (Antoniadis et al., 2016; Linares et al., 2018; Strader et al., 2019; Linares, 2020) and fastest spinning NS. The fastest spinning pulsar known so far spins at a frequency of 716 Hz (PSR J1748-2446ad spinning at about 1.4 ms; Hessels et al., 2006) in the globular cluster Terzan 5. Various reasons have been discussed to explain the lack of pulsars spinning faster than 1 ms, for example an additional torque (from gravitational wave emission) that acts at high frequencies and spins down the pulsar (Papaloizou & Pringle, 1978; Wagoner, 1984; Bildsten, 1998; Bhattacharyya & Chakrabarty, 2017). Since pulsar spins depend on the accretion phase, we investigate here the fastest spins possible for pulsars. Despite relatively large mass uncertainties, NSs in spider binaries tend towards higher masses (median $M_{\rm NS}\sim 1.8\,{\rm M}_{\odot}$; Linares, 2020) in comparison to other systems like double NSs or wide NS+white dwarf (WD) binaries (Demorest et al., 2010; Antoniadis et al., 2013; Linares, 2020).
@@
In this study, we carry out detailed binary evolution simulations, calculating the evolution of the pulsar spin and taking into account isolated and mass-transfer evolution. The main aim is to investigate the effect of various assumptions of binary interaction on the pulsar spin evolution and to constrain these when comparing with observed spider properties. In Section 2, we describe the initial parameters of the simulated tracks, the added physics of pulsar spin evolution, and the treatment of the pulsar wind irradiation. Section 3 presents our simulated grids in the context of the observed population of spiders. In Section 4, we compare our results with those in the literature and discuss the caveats involved, followed by the conclusions of our study in Section 5.
@@ 2.1 Initial pulsar properties and binary physics
We use the detailed stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA; Paxton et al., 2011, 2013, 2015, 2018, 2019) to carry out our calculations for the evolution of spiders and included pulsar spin evolution. One of the benefits of using a detailed stellar code like MESA  is that compared to rapid population synthesis codes, detailed codes provide a more self-consistent calculation of the mass-transfer phase along with a physically motivated estimation of the stability of binary interaction. All the initial binaries in this study have a low- to intermediate-mass main-sequence star (MS; range of 1.0 to 4.5 M) and a point mass NS (range of 1.3 to 2.0 M). The initial orbital periods range from 0.6 to 4 d. The simulations are run at Solar metallicity (0.0142; Asplund et al., 2009). Since most of the observed spider binaries with dynamical mass measurements are within the Milky Way (Linares, 2020), we are focusing on Galactic spiders for our study. The physics in the MESA  grids follows the description in Fragos et al. (2023). We treat our numerically failed runs similar to Fragos et al. (2022), by identifying failed binaries and re-running them with changed parameters that will help computational stability and reach an acceptable termination condition. We discuss our treatment of the numerically failed binaries in Appendix A.
@@ 2.2 Isolated pulsar phase
As initial pulsar spin parameters, we take a spin of $1\,\rm s$ and a spin period derivative of $10^{-15}\,\rm s/s$. The corresponding magnetic field strength is $10^{12}\,\rm G$. These values correspond to the average spin parameters of observed young pulsars. We model the pulsar spin evolution by identifying different stages of the evolution and applying the corresponding equations, as described in the following sections. Pulsars in detached binaries, where the binary components are orbiting each other without transferring mass, are essentially assumed to spin down as isolated pulsars as their spins and magnetic fields are not affected by the companion. Since pulsars are rapidly spinning dipoles, they lose energy as they spin down in the form of magnetic dipole radiation, provided the dipole has an inclination. Hence, the spin-down of a pulsar results in the following decrease in its angular frequency ($\dot{\Omega}$),
@@
where $\mu$ is the pulsar magnetic moment ($\mu\approx BR^{3}$), $\alpha$ is the angle between the rotation axis and magnetic axis, $\mu_{0}$ is the permeability of free space (it is unity in CGS units), $I$ is the pulsar moment of inertia (approximated as $52M_{\rm NS}R_{\rm NS}^{2}/$ 5 for a solid sphere, where $M_{\rm NS}$ and $R_{\rm NS}$ are the NS mass and radius), and $c$ is the speed of light. We assumed $\alpha=45\degree$, which is a simplistic approximation, as the inclination would decay with time (Spitkovsky, 2006; Johnston & Karastergiou, 2017). The full range of $\alpha$ values would give an order of magnitude difference in $\dot{\Omega}$, at most (Spitkovsky, 2006), hence, we do not vary this parameter in our study and fix it at an intermediate value.
We evolve the isolated pulsar surface magnetic field ($B$) decay following Chattopadhyay et al. (2020),
@@
where $B^{i}$ is the initial surface magnetic field, $B_{\rm min}$ is the minimum magnetic field (we take $B_{\rm min}=10^{8}\,\rm G$; Zhang & Kojima, 2006; Osłowski et al., 2011), $t$ is time, and $\tau_{\rm d}$ is the magnetic field decay timescale (we take $\tau_{\rm d}=3\,\rm Gyr$; Ye et al., 2019). The superscripts $i$ and $f$ stand for initial and final values. Previous studies have explored $\tau_{\rm d}$ values ranging from a few Myr to a few Gyr, with decay timescales $\gtrsim 2\,\rm\,Gyr$ being favoured (Kiel et al., 2008; Ye et al., 2019; Chattopadhyay et al., 2020).
@@ 2.3 Interacting binary phase
Once the companion to the pulsar fills its Roche lobe and the mass-transfer phase begins, accreted matter will carry angular momentum and spin-up the pulsar. We consider mass accretion when there is a persistent accretion disc and use the criteria by Dubus et al. (1999) that describes the limiting mass-transfer rate below which thermal-viscous instabilities would prevent accretion. This limit is as follows,
@@
The change in the NS mass ($\dot{M}_{\rm NS}$) is determined by the efficiency of accretion ($\epsilon$) which is assumed to be equal to $1-\beta$, where $\beta$ is the fraction of mass transferred by the companion that is lost from the vicinity of the NS as a fast isotropic wind, taking away angular momentum from the NS. The rest of the material ($1-\beta$) is accreted by the NS, until the mass-transfer rate reaches the Eddington limit, beyond which all accretion is limited. Since the term $\beta$ affects the amount of accretion, lower values of $\beta$ result in faster spin-up of the pulsar compared to higher values. Additionally, the value of $\beta$ affects the orbital evolution following the equation for the relative change in orbital separation ($\dot{a}/a$; Tauris & van den Heuvel, 2006), assuming other changes in orbital angular momentum are represented by $\dot{J}_{\rm orb}$,
@@
Theoretical and observational studies of various types of binaries have shown that the parameter $\beta$ could be more than 0.5 in LMXBs despite the evolution being mostly sub-Eddington (Tauris & Savonije, 1999; Ritter & King, 2001; Podsiadlowski et al., 2002; Jacoby et al., 2005; Zhang et al., 2011; Antoniadis et al., 2012; Tauris et al., 2012). Some possible reasons for this are the propeller effect (Illarionov & Sunyaev, 1975) and thermal-viscous instabilities in the accretion disc (Pringle, 1981; Mineshige, 1993; van Paradijs, 1996); all effects that would prevent accretion. Studies of accreting pulsars use values in the range of 0.5 to 0.7 (Podsiadlowski et al., 2002; Chen et al., 2013; Jia & Li, 2014, 2015; Misra et al., 2020; Wang et al., 2021). We add a prescription for thermal-viscous accretion disc instabilities (explained further below in Section 2.3) and explore three values of $\beta$ (0.0, 0.3, and 0.7) to study its effect on the pulsar spin evolution. This fractional value $\beta$ is also a proxy for the change in angular momentum when mass is accreted. We include $\beta$ even though we have a separate treatment for the propeller phase (discussed below in Section 2.3.2) and the aforementioned disc instabilities, since these have large uncertainties and there might still be inefficient transfer of angular momentum due to some unforeseen physical process.
The accreted matter suppresses the magnetic field due to rapid Ohmic decay (Konar & Bhattacharya, 1997, 1999a, 1999b), which we include as an exponential decay with accreted mass (Osłowski et al., 2011; Chattopadhyay et al., 2020), as follows,
@@
In the above equation, $\Delta M_{\rm NS}$ is the accreted matter and $M_{\rm d}$ is the magnetic field mass decay scale (0.025, although other values for $M_{\rm d}$ have been explored in the range of 0.0033 to 0.05; Osłowski et al., 2011; Chattopadhyay et al., 2020). The accreted matter also causes a change in the angular momentum of the NS which can be estimated as,
@@
where $\kappa$ is an efficiency factor and $\omega_{\rm diff}$ is the angular velocity of the incoming material (Kiel et al., 2008; Chattopadhyay et al., 2020) that is defined by,
@@
In the above equation, $r_{\rm cor}$ is the co-rotation radius, which is defined as $r_{\rm cor}=(GM_{\rm NS}/\Omega^{2})^{1/3}$ and $r_{\rm mag}$ is the magnetospheric radius, which is defined as follows (Kiel et al., 2008; Chattopadhyay et al., 2020),

### v2

@@ Abstract
Context. Compact binary millisecond pulsars (MSPs) with orbital periods $\lesssim 1$d are key to understanding binary evolution involving massive neutron stars (NSs). Due to the ablation of the companion by the rapidly spinning pulsar, these systems are also known as spiders and categorized into two main branches: redbacks (RBs; companion mass in the range of 0.1 to 0.5 ${\rm M}_{\odot}$) and black widows (BWs; companion mass $\lesssim$ 0.1 ${\rm M}_{\odot}$).
Aims. We present models of low- and intermediate-mass X-ray binaries and compare them with observations of Galactic spiders (including the presence or absence of hydrogen lines in their optical spectra), and we constrain and quantify the interaction between the pulsar and the companion.
Methods. Using MESA, we created the allowed initial parameter space. For the first time in MESA , we also included the detailed evolution of the pulsar spin and modeled the irradiation of the companion by the pulsar wind.
Results. Efficient mass accretion onto the NS (i.e., at least $70\%$ of the mass transferred is accreted) with an X-ray irradiated disk followed by strong irradiation of the companion can explain most of the properties of the observed spiders. Our RB evolutionary tracks continue to the BW regime, connecting the two branches of spiders. Our models explain the lack of hydrogen in some observed BWs with ultra-light companions. During accretion induced spin up, the mass required to spin up an NS to sub-milliseconds is high enough to collapse it into a black hole. Finally, after analyzing the formation of RB-like spiders with giant companions and orbital periods of several days (huntsmen), we conclude that they are unlikely to produce super-massive NSs (maximum accreted mass $\lesssim$0.5 ${\rm M}_{\odot}$).
@@ 1 Introduction
The formation of radio millisecond pulsars (MSPs) has long been connected to low-mass X-ray binaries (LMXBs; (Alpar et al., 1982; Radhakrishnan & Srinivasan, 1982; Bhattacharya & van den Heuvel, 1991)). Observed MSPs have short spin periods ($\lesssim 30$ ms) and low surface magnetic fields ($\sim 10^{8}$ G). The neutron stars (NSs) in LMXBs accrete matter that spins them up to millisecond periods (Alpar et al., 1982). This phase is observed as an X-ray phase and is also called the recycling phase due to the pulsar spin-up. Observations of accreting X-ray MSPs confirmed the link between MSPs and accreting LMXBs (Wijnands & van der Klis, 1998; Archibald et al., 2009). The NS surface magnetic field has also been suggested to decrease with the accreted mass due to Ohmic decay, explaining the observed magnetic field strengths of MSPs (Konar & Bhattacharya, 1997; Zhang, 1998; Konar & Bhattacharya, 1999a, b; Cumming et al., 2001). Compact binary MSPs are a sub-category of MSP binaries with compact orbits and eclipses in their radio signals. In compact binary MSPs, after the mass-transfer phase or Roche-lobe overflow (RLO) has ended, the highly spinning pulsar loses energy in the form of energetic particles that irradiate the outer surface of the companion, leading to mass loss (Archibald et al., 2009). Initial observations of eclipsing MSPs showed that some of them have very low mass companions (0.02 to 0.05 ${\rm M}_{\odot}$; Fruchter et al., 1988; Stappers et al., 1996a, 1999), which could be explained by mass loss due to pulsar wind irradiation. Evidence of irradiation-induced mass loss has since been observed in multiple eclipsing compact MSPs (Kluzniak et al., 1988; Ruderman et al., 1989a, b; Stappers et al., 1996b; Polzin et al., 2020; Kumari et al., 2024).
@@
Due to the cannibalistic nature of this interaction, these compact binary MSPs, with observed orbital periods $P_{\rm orb}\lesssim 1\,\rm d$ (Pallanca et al., 2012; Romani et al., 2012; Kaplan et al., 2013; Breton et al., 2013; Roberts, 2013), are also called spider pulsars. So far, about 60 spiders have been observed and studied in the Galaxy (for instance, see Abdo et al., 2009; Pletsch et al., 2012; Pallanca et al., 2012; Romani et al., 2012; Kaplan et al., 2013; Breton et al., 2013). There are two main categories of spiders (Roberts, 2013) based on the masses of the companions to the pulsar (denoted as $M_{\rm c}$): redbacks (RBs), which have companions with masses $0.1{\,{\rm M}_{\odot}}\lesssim M_{\rm c}\lesssim 0.5\,{\rm M}_{\odot}$, and black widows (BWs), whose companions have masses of $M_{\rm c}<0.1\,{\rm M}_{\odot}$. The orbital period distribution is similar for both types ($\lesssim$1 d).
@@
Following classical binary evolution, the observed orbital periods of all spiders have not been explained, as they fall short by about an order of magnitude (Podsiadlowski et al., 2002). Fast winds that leave a binary isotropically and without interacting with anything else have a widening effect on the binary orbit (Soberman et al., 1997). Therefore, the involvement of this energetic radio pulsar wind has been invoked to explain the orbital properties of recycled pulsars (Ergma & Fedorova, 1992; Podsiadlowski et al., 2002; Benvenuto et al., 2014, 2015; Chen et al., 2013). There are a few ideas that have been proposed to explain spider MSPs. Chen et al. (2013) proposed strong and weak irradiation producing RBs and BWs, respectively. Another group studied the formation of spiders by including X-ray irradiation feedback, and they found that in order to reproduce the orbital parameters of spiders, cyclic mass transfer due to the effect of X-rays is required (Benvenuto et al., 2014, 2015). Ginzburg & Quataert (2020) and Ginzburg & Quataert (2021) argue that sustained magnetic braking (which couples to stellar wind and leads to angular momentum loss) is needed to reproduce spiders and that ablation alone is not sufficient. Some spiders are also thought to come from ultra-compact X-ray binaries (UCXBs), which are a sub-category of LMXBs with hydrogen-deficient companions and orbital periods of less than an hour (see review by Nelemans & Jonker, 2010).
@@
Since spider binaries have passed through a long mass-transfer phase, it has been suggested that they host some of the most massive (Antoniadis et al., 2016; Linares et al., 2018; Strader et al., 2019; Linares, 2020) and fastest spinning NSs. The fastest spinning pulsar known so far spins at a frequency of 716 Hz (PSR J1748-2446ad spinning at about 1.4 ms; Hessels et al., 2006) in the globular cluster Terzan 5. Various reasons have been discussed to explain the lack of pulsars spinning faster than 1 ms, for example an additional torque (from gravitational wave emission) that acts at high frequencies and spins down the pulsar (Papaloizou & Pringle, 1978; Wagoner, 1984; Bildsten, 1998; Bhattacharyya & Chakrabarty, 2017). Since pulsar spins depend on the accretion phase, in this work we investigate the fastest spins possible for pulsars. Despite relatively large mass uncertainties, NSs in spider binaries tend toward higher masses (median $M_{\rm NS}\sim 1.8\,{\rm M}_{\odot}$; Linares, 2020) in comparison to other systems, such as double NSs or wide NS plus white dwarf (WD) binaries (Demorest et al., 2010; Antoniadis et al., 2013; Linares, 2020).
@@
In this study, we carry out detailed binary evolution simulations, calculating the evolution of the pulsar spin and taking into account isolated and mass-transfer evolution. The main aim is to investigate the effect of various assumptions of binary interaction on the pulsar spin evolution and to constrain them when comparing with observed spider properties. In Section 2, we describe the initial parameters of the simulated tracks, the added physics of pulsar spin evolution, and the treatment of the pulsar wind irradiation. Section 3 presents our simulated grids in the context of the observed population of spiders. In Section 4, we compare our results with those in the literature and discuss the caveats involved. This is followed by the conclusions of our study in Section 5.
@@ 2.1 Initial pulsar properties and binary physics
We use the detailed stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA; Paxton et al., 2011, 2013, 2015, 2018, 2019) to carry out our calculations for the evolution of spiders and included pulsar spin evolution. One of the benefits of using a detailed stellar code like MESA  is that compared to rapid population synthesis codes, detailed codes provide a more self-consistent calculation of the mass-transfer phase along with a physically motivated estimation of the stability of binary interaction. All the initial binaries in this study have a low- to intermediate-mass main-sequence star (MS; range of 1.0 to 4.5 ${\rm M}_{\odot}$) and a point mass NS (range of 1.3 to 2.0 ${\rm M}_{\odot}$). The initial orbital periods range from 0.6 to 4 d. The simulations are run at Solar metallicity (0.0142; Asplund et al., 2009). Since most of the observed spider binaries with dynamical mass measurements are within the Milky Way (Linares, 2020), we are focusing on Galactic spiders for our study. The physics in the MESA  grids follows the description in Fragos et al. (2023). We treat our numerically failed runs similar to Fragos et al. (2022), by identifying failed binaries and rerunning them with changed parameters that will help computational stability and reach an acceptable termination condition. We discuss our treatment of the numerically failed binaries in Appendix A.
@@ 2.2 Isolated pulsar phase
As initial pulsar spin parameters, we take a spin of $1\,\rm s$ and a spin period derivative of $10^{-15}\,\rm s/s$. The corresponding magnetic field strength is $10^{12}\,\rm G$. These values correspond to the average spin parameters of observed young pulsars. We model the pulsar spin evolution by identifying different stages of the evolution and applying the corresponding equations, as described in the following sections. Pulsars in detached binaries, where the binary components are orbiting each other without transferring mass, are essentially assumed to spin down as isolated pulsars as their spins and magnetic fields are not affected by the companion. Since pulsars are rapidly spinning dipoles, they lose energy as they spin down in the form of magnetic dipole radiation, provided the dipole has an inclination. Hence, the spin-down of a pulsar results in the following decrease in its angular frequency ($\dot{\Omega}$):
@@
where $\mu$ is the pulsar magnetic moment ($\mu\approx BR^{3}$), $\alpha$ is the angle between the rotation axis and magnetic axis, $\mu_{0}$ is the permeability of free space (it is unity in CGS units), $I$ is the pulsar moment of inertia (approximated as $52M_{\rm NS}R_{\rm NS}^{2}/$ 5 for a solid sphere, where $M_{\rm NS}$ and $R_{\rm NS}$ are the NS mass and radius), and $c$ is the speed of light. We assumed $\alpha=45\degree$, which is a simplistic approximation, as the inclination would decay with time (Spitkovsky, 2006; Johnston & Karastergiou, 2017). The full range of $\alpha$ values would give an order of magnitude difference in $\dot{\Omega}$, at most (Spitkovsky, 2006); hence, we did not vary this parameter in our study and fixed it at an intermediate value.
We evolved the isolated pulsar surface magnetic field ($B$) decay following Chattopadhyay et al. (2020):
@@
where $B^{i}$ is the initial surface magnetic field, $B_{\rm min}$ is the minimum magnetic field (we take $B_{\rm min}=10^{8}\,\rm G$; Zhang & Kojima, 2006; Osłowski et al., 2011), $t$ is time, and $\tau_{\rm d}$ is the magnetic field decay timescale (we take $\tau_{\rm d}=3\,\rm Gyr$; Ye et al., 2019). The superscripts $i$ and $f$ stand for initial and final values. Previous studies have explored $\tau_{\rm d}$ values ranging from a few Myr to a few Gyr, with decay timescales $\gtrsim 1\,\rm\,Gyr$ being favored (Kiel et al., 2008; Ye et al., 2019; Chattopadhyay et al., 2020).
@@ 2.3 Interacting binary phase
Once the companion to the pulsar fills its Roche lobe and the mass-transfer phase begins, accreted matter will carry angular momentum and spin up the pulsar. We consider mass accretion when there is a persistent accretion disk and use the criteria by Dubus et al. (1999) that describes the limiting mass-transfer rate below which thermal-viscous instabilities would prevent accretion. This limit is as follows:
@@
The change in the NS mass ($\dot{M}_{\rm NS}$) is determined by the efficiency of accretion ($\epsilon$), which is assumed to be equal to $1-\beta$, where $\beta$ is the fraction of mass transferred by the companion that is lost from the vicinity of the NS as a fast isotropic wind, taking away angular momentum from the NS. The rest of the material ($1-\beta$) is accreted by the NS, until the mass-transfer rate reaches the Eddington limit, beyond which all accretion is limited. Since the term $\beta$ affects the amount of accretion, lower values of $\beta$ result in a faster spin-up of the pulsar compared to higher values. Additionally, the value of $\beta$ affects the orbital evolution following the equation for the relative change in orbital separation ($\dot{a}/a$; Tauris & van den Heuvel, 2006), assuming other changes in orbital angular momentum are represented by $\dot{J}_{\rm orb}$:
@@
Theoretical and observational studies of various types of binaries have shown that the parameter $\beta$ could be more than 0.5 in LMXBs despite the evolution being mostly sub-Eddington (Tauris & Savonije, 1999; Ritter & King, 2001; Podsiadlowski et al., 2002; Jacoby et al., 2005; Zhang et al., 2011; Antoniadis et al., 2012; Tauris et al., 2012). Some possible reasons for this are the propeller effect (Illarionov & Sunyaev, 1975) and thermal-viscous instabilities in the accretion disk (Pringle, 1981; Mineshige, 1993; van Paradijs, 1996); all effects that would prevent accretion. Studies of accreting pulsars use values in the range of 0.5 to 0.7 (Podsiadlowski et al., 2002; Chen et al., 2013; Jia & Li, 2014, 2015; Misra et al., 2020; Wang et al., 2021). We add a prescription for thermal-viscous accretion disk instabilities (explained further below in Section 2.3) and explore three values of $\beta$ (0.0, 0.3, and 0.7) to study its effect on the pulsar spin evolution. This fractional value $\beta$ is also a proxy for the change in angular momentum when mass is accreted. We include $\beta$ even though we have a separate treatment for the propeller phase (discussed below in Section 2.3.2) and the aforementioned disk instabilities, since these have large uncertainties and there might still be inefficient transfer of angular momentum due to some unforeseen physical process.
The accreted matter suppresses the magnetic field due to rapid Ohmic decay (Konar & Bhattacharya, 1997, 1999a, 1999b), which we include as an exponential decay with accreted mass (Osłowski et al., 2011; Chattopadhyay et al., 2020), as follows:
@@
In the above equation, $\Delta M_{\rm NS}$ is the accreted matter and $M_{\rm d}$ is the magnetic field mass decay scale (0.025, although other values for $M_{\rm d}$ have been explored in the range of 0.0033 to 0.05; Osłowski et al., 2011; Chattopadhyay et al., 2020). The accreted matter also causes a change in the angular momentum of the NS, which can be estimated as
@@
where $\kappa$ is an efficiency factor and $\omega_{\rm diff}$ is the angular velocity of the incoming material (Kiel et al., 2008; Chattopadhyay et al., 2020) that is defined by
@@
In the above equation, $r_{\rm cor}$ is the co-rotation radius, which is defined as $r_{\rm cor}=(GM_{\rm NS}/\Omega^{2})^{1/3}$ and $r_{\rm mag}$ is the magnetospheric radius, which is defined as follows (Kiel et al., 2008; Chattopadhyay et al., 2020):